# Teaching an Old Dog New Tricks: Verifiable FHE Using Commodity Hardware

### v1

@@ 7.2. Malicious Private Set Intersection
Most efficient PSI protocols do not achieve (full) malicious security (Cong et al., 2021; Chen et al., 2017b, 2018; Kiss et al., 2017). Some constructions do (Nevo et al., 2021; Rosulek and Trieu, 2021; Pinkas et al., 2020; De Cristofaro et al., 2010; Raghuraman and Rindal, 2022; Rindal and Schoppmann, 2021; Sun et al., 2024) but at the price of high offline communication cost (see section 9.5). Using a state-of-the-art semi-honest PSI (Cong et al., 2021) scheme based on FHE, we show how Argos can be extended to make the scheme fully secure against a malicious server. Here, the main difference from the PIR setup is that the server inputs are not public, but private. State-of-the-art FHE schemes are not circuit private. That means the client might be able to learn information regarding the server’s private input by looking at the noise contained in the output. As it is standard in FHE-based PSI schemes, this requires to mask the server’s input using oblivious pseudorandom functions (OPRFs) to provide privacy for the server even in the presence of malicious clients. Moreover, security of the FHE scheme is enough to also provide privacy for the client, even in the presence of a “malicious” server. This is because the ideal PSI functionality considered in these works assumes a limited attacker with no observability of the client reactions (Chen et al., 2018). However, this might not hold in a real-world deployment, as we explained before.
Verifiable PSI. Once deployed in a real-world setting, the attacker might be able to observe some of the client’s reactions. Attacks on correctness might have serious security implications, sometimes translating to attacks on privacy (see example for private contact discovery in section 3.3). Fortunately, circuit-level integrity—as provided by Argos—can guarantee verifiable PSI, which means that verification of the attestation transcript succeeds if and only if the server correctly executed the PSI circuit on a private set. Nevertheless, for full malicious security, we still need to ensure that the server evaluated the over the correct set.
@@
Similarly to Authenticated PIR, Argos can be extended to include a commitment to the dataset (e.g. Merkle tree or hash-based commitment scheme) in the attestation transcript. Here too, a public commitment needs to be made available out-of-band to the client, through a transparency log, for instance. Upon receiving a response to its request, the client first verifies the transcript and aborts if it does not match the correct commitment. Formally, the extensions we made for PIR are almost sufficient, we just need to tweak the wrapper function over to match our PSI functionality. Instead of only checking that are well formed, the wrapping function should compute the corresponding OPRF values for . The server’s privacy is guaranteed by the OPRF, the soundness of the attestation scheme, and the fact that commitment schemes are hiding i.e., they do not leak information about the server’s set.
@@ 8.1. Security Monitor
Our security monitor is based on the X86 version of Tyche (Castes et al., 2023), a Rust micro-hypervisor that exposes hardware resources and isolation primitives to better protect trust domains. Note that other “small” hypervisors such as XEN (Barham et al., 2003) are at least one order of magnitude larger or closed source and micro-kernels such as SeL4 (Klein et al., 2009). We implement our attestation scheme on top of Tyche, including support for the hardware TPM and the interface to manage attested transcripts. We are also able to simplify some aspects of the platform, for instance, by removing side channel protections from Tyche (e.g., cache and CPU flush). One of the implementation challenges was to adapt Tyche and the Linux driver to support large (i.e., several GBs) enclaves with our custom runtime and manage memory allocation accordingly. To enforce the integrity of the security monitor, we use the TPM and a dynamic root of trust ((e.g., Intel TXT or AMD DRMT (Futral and Greene, 2013; Nie, 2007))) to implement standard measure boot (see primer in appendix A). At the end of this process, two TPM registers (PCRs) are set with hash values that uniquely identify the security monitor binary. Later, when the security monitor needs to attest an application transcript, it will load the transcript hash in a third PCR and request a TPM Quote (a signature over select PCRs).
@@ 8.3. Custom Runtime
The execution environments provided by Argos are similar to bare-metal environments with no OS support. That means no syscalls, no program loader, no memory allocation, nor access to shared libraries. A runtime is required to support programs and modern libraries like SEAL that leverage many services from the OS. Gramine (Tsai et al., 2017) is such a runtime and already supported by Tyche. It was conceived to securely run off-the-shelve applications on Intel SGX. As a result, it implements all Linux syscalls and also allow extensive support for shared library and memory management. However, because of its versatility, it is rather large (20KLOC) and adds some non-negligible overhead (Project, 2024b). In order to reduce the TCB and improve performance, we implement our own custom runtime. It is minimal (870LOC) and consists of a simple memory allocator (Peterson and Norman, 1977) and a handful of syscall handlers required by our application. For instance, we implement our own open and read syscalls to supply randomness from RDRAND when an application attempts to read /dev/urandom. We write our runtime to interface with MUSL (musl, 2024) by interjecting on system calls and providing our own handlers. That means all applications and libraries (including SEAL (Chen et al., 2017a)) need to be compiled statically and linked using our MUSL library, which requires some engineering efforts to port applications but makes it possible to obtain high performance and a very light runtime.
@@ 9.1. Testbench
We prototype Argos on a local Dell Optiplex machine from 2017 equipped with an Intel Core i7-7700 processor clocked at 3.6GHz, a discrete TPM 2.0, and 8GB of RAM. We run our security monitor (adapted from Tyche (Castes et al., 2023) see section 8.1) directly on the hardware and evaluate our benchmarks using our custom SEAL runtime or Gramine (Tsai et al., 2017). We run Ubuntu 22.04 with the Linux kernel v6.2 and SEAL 4.1 with the BFV scheme. All benchmarks are all single-threaded, compiled using clang++ and using O3 optimization level. We also compare Argos with two other TEE platforms that have the following specs. SGXv2 running on an Azure server equipped with an Intel Xeon Platinum 8370C CPU with a frequency of 2.80GHz. AWS Nitro Enclaves (AWS, 2024a) running on an AWS server equipped with an Intel Xeon Platinum 8259CL CPU with a frequency of 2.50GHz. The SGXv2 setup uses Gramine as a runtime while Nitro Enclave uses a docker container environment that runs its own Linux kernel. Note that under our threat model, both SGXv2 and AWS Nitro Enclaves are insecure and are only used as comparison points for performance. See table 7 for a more complete comparison. All numbers are obtained by averaging values over 10 runs with minimal variance.
@@ 9.2. TCB Evaluation
Detailed count for lines of code (LOCs) can be found in 3. Our TCB is small with a minimum of 40KLOC for our PIR application and a total of 60KLOC for our PSI application. Note that we did not account for the microcode used it the DRoT or the TPM as it is not public information, and we consider these elements as part of the hardware. This should be compared to other TEEs and system projects that try to keep a small footprint. For instance, Nitro Enclaves include an entire Linux kernel which blows up their TCB in the range of MLOC, several orders of magnitude bigger than us. On the other hand, SGX would not include the security monitor (18KLOC), but relies on a significant amount of microcode. As a point of comparison, XEN (Barham et al., 2003) (a popular mini-hypervisor), an Coreboot (Borisov, 2009) (a minimal BIOS) each total 200KLOC.
@@ 9.3. Microbenchmarks
The vTPM signature requires a vmcall and a context switch to the security monitor, which incurs a small overhead (about 100us) compared to the application-level signature.
@@
However, they are only 5x faster than discrete TPM at best.
@@
Enclave Fixed Costs. We evaluate the different fixed costs of different operations over the lifetime of an enclave. To obtain measurements, we instrument a dummy enclave for which we allocate 1GB of memory. Setup time is measured between the time a command line program requests the operating system to start an enclave and when execution reaches the main function of the enclave application. This means it includes the allocation of enclave resources, loading of the binary, page table setup, binary measurement, and runtime initialization. We can see that Nitro Enclaves have the highest setup time mostly due to the fact that they contain an entire linux kernel that needs to be booted. SGXv2 also presents some large startup time. This is mostly due to all the special instructions required to start an enclave and the time it takes Gramine to initialize. As indicated in Gramine’s documentation, initialization can initiate more than 200 OCALL which all incur expensive context switches, with architectural flush (CPU caches and state) that cost 8,000 — 12,000 cycles each (Project, 2024b). This overhead can also be observed when running Argos with Gramine. In comparison, our custom runtime is initialized with minimal overhead, which might be essential for latency-sensitive applications.
@@
For Nitro Enclaves, it requires to wait for Linux to shutdown, which aslo adds delay.
@@
Transcript Size. Our transcript is quite small and composed of 2 PCR values used by the DRoT, the initial value of some key registers for the execution environment, the hash of our binary, the hash of messages exchanged by the application, the hash of some server input if needed. It also contains the signature and the TPM public key along with the TPM certificates endorsed by the manufacturer. Concretelly, we measure a small transcript size of only 1,407 bits, negligible in front of the size of an FHE ciphertext (500KB).
@@ 9.4. FHE Evaluation
We evaluate the overhead of circuit-level vFHE in Argos compared to semi-honest FHE. Results can be seen in table 6. We use the same benchmarks than previous work on vFHE (Viand et al., 2023) and compile and evaluate the different circuits using the SEAL library (Chen et al., 2017a). The results from Viand. et al. highlight the impractical performance overheads when using cryptographic proofs to build vFHE: between 6 and 7 orders of magnitude, with the exception of Pinocchio for the Tiny benchmark. Their baseline is also relatively slow, hinting that they might have disable hardware acceleration. Their implementation of FHE-in-TEE also shows poor performance probably due to the limitation of SGXv1 (limited 128MB EPC memory and unreliable timers). For a fair comparison, we run the benchmarks of SGXv2 and Nitro Enclaves. We compiled the benchmarks on one machine and used the same binary across or different setups. Here, the performance looks much better with single-digit overheads and even small speedups for Nitro Enclaves. Indeed, once setup, running a program in an enclave is equivalent to running it with bare-metal performance. We tried to keep the runtime as similar as possible between the different configurations, but small differences might explain these results. For Argos, we evaluate two different configurations. One using the Gramine runtime (Argos+G) and one using our custom runtime. Our custom runtime performs better than Gramine. This might be for a couple reasons including that the MUSL library tends to be faster than Glibc, our custom runtime simply ignores interrupts sent by the OS scheduler (the enclave is not de-scheduled), and our custom syscall handlers handle less corner cases and are more efficient (for instance, to allocate memory). In conclusion, Argos shows performance comparable to commercial TEE while offering better security. Our custom runtime makes it possible to reduce the TCB and improve performance at the cost of some engineering effort to port applications.
@@ 9.5. Authenticated Private Information Retrieval
We evaluate the overhead of authenticated PIR using Argos compared to semi-honest unauthenticated PSI. We use the SealPIR library from Angle et. al. (Angel et al., 2018). We use the BFV scheme with polynomials of degree 4096. The plaintext modulus is set to 20. We evaluate performance for a database of size where each element measures 128 bytes. We allocate 4GB of static memory to our enclave. We instrument the code to hash the database, queries, and responses and to attest the final transcript once all queries have been processed. Results can be found in table 6. We compare two configurations for Argos. One using the Gramine runtime (Argos+G) and one using our custom runtime. For Argos+G, the server setup –which is a fixed cost– shows an overhead of mostly due to overhead setting up the enclave (+55% alone) and hashing the database (+16%). For query processing, we measure an overhead of for fetching one element. Here, hashing the response before sending it to the client takes a negligible amount of time. Taking into account the signing of the final transcript, the overhead goes up to for retrieving one element. Batching queries here would amortize costs down to a minimum of for numerous queries. For Argos using our custom runtime, overhead goes down to 21% for setup (mostly due to hashing of the batabase) and 1% for query processing. This is due to our more efficient handling of system calls. The only cost left is attestation that can be amortized by batching requests (see section 6.2). Once again, our custom runtime offers better performance than Gramine. For all setups, Argos’s transcript incurs negligible (¿1%) communication overhead.
Recent work on single-server authenticated PIR using cryptographic constructions by Colombo. et al., while providing better security guarantees, show a performance overhead of 30-100x (Colombo et al., 2023). Other recent work, VeriSimplePIR (de Castro and Lee, 2024), show more reasonable online overheads (13%-20%) but incur significant offline communication in the same order of magnitude than the database itself (2GiB communication for a 4GiB dataset). This also makes updating the database extremely difficult, while our scheme only requires updating a Merkle tree and communicating the commitment.
@@ 9.6. Authenticated Private Set Intersection
Attested Transcript. Our remote attestation scheme makes it possible to amortize the cost of attestation over messages and queries. In the PSI scheme we use, each query requires two rounds of communication (one round for OPRF, and one for the query itself), plus one initial attestation, which would bring the overhead for one query to . Instead, we only need to pay the cost of one attestation operation per batch. The PSI scheme we use already supports batching but only up to a certain client set size (this is a limitation of FHE-based PSI schemes). For our given server set size, we empirically measure a non-negligible failure rate for client set size above elements. A client who would want to compute the intersection of a bigger set would have to do so using several queries. That means that the maximum overhead of for a client set size of 1 can be amortized for larger client set sizes. For instance, a client set size of , computing the complete intersection would take 10 queries and overhead would be reduced back to . This phenomenon would be heightened for more efficient PSI schemes as the cost of TPM signing is constant.

### v2

@@ 7.2. Malicious Private Set Intersection
Most efficient PSI protocols do not achieve (full) malicious security (Cong et al., 2021; Chen et al., 2017b, 2018; Kiss et al., 2017). Some constructions do (Nevo et al., 2021; Rosulek and Trieu, 2021; Pinkas et al., 2020; De Cristofaro et al., 2010; Raghuraman and Rindal, 2022; Rindal and Schoppmann, 2021; Sun et al., 2024) but at the price of high offline communication cost (see Section 9.5). Using a state-of-the-art semi-honest PSI (Cong et al., 2021) scheme based on FHE, we show how Argos can be extended to make the scheme fully secure against a malicious server. Here, the main difference from the PIR setup is that the server inputs are not public, but private. State-of-the-art FHE schemes are not circuit private. That means that the client might be able to learn information regarding the server’s private input by looking at the noise contained in the output. FHE-based PSI schemes mask the server’s input using oblivious pseudo random functions (OPRFs) to provide privacy for the server even in the presence of malicious clients. Moreover, the security of the FHE scheme is argued to also provide privacy for the client, even in the presence of a “malicious” server. This is because the ideal PSI functionality considered in these works assumes a limited attacker with no observability of the client reactions (Chen et al., 2018). However, this might not hold in a real-world deployment, as we explained before.
Verifiable PSI. Once deployed in a real-world setting, the attacker might be able to observe some of the client’s reactions. Attacks on correctness might have serious security implications, sometimes translating to attacks on privacy (see example of private contact discovery in Section 3.3). Fortunately, circuit-level integrity—as provided by Argos—can guarantee verifiable PSI, which means that verification of the attestation transcript succeeds if and only if the server correctly executed the PSI circuit on a private set. Nevertheless, for full malicious security, we still need to ensure that the server evaluates PSI over the correct set.
@@
Similarly to Authenticated PIR, Argos can be extended to include a commitment to the dataset (e.g., a Merkle tree or hash-based commitment scheme) in the attestation transcript. Here too, a public commitment needs to be made available out-of-band to the client, through a transparency log, for instance. Upon receiving a response to its request, the client first verifies the transcript and aborts if it does not match the correct commitment. Formally, the extensions we made for PIR are almost sufficient, we just need to tweak the wrapper function over to match our PSI functionality. Instead of only checking that is well formed, the wrapping function should compute the corresponding OPRF values for . The server’s privacy is guaranteed by the OPRF, the soundness of the attestation scheme, and the fact that commitment schemes are hiding i.e., they do not leak information about the server’s set. The security argument can be found in Appendix F.
@@ 8.1. Security Monitor
Our security monitor is based on the X86 version of Tyche (Castes et al., 2023), a Rust micro-hypervisor that exposes hardware resources and isolation primitives to better protect trust domains. Note that other “small” hypervisors such as XEN (Barham et al., 2003) are at least one order of magnitude larger or closed source. We implement our attestation scheme on top of Tyche, including support for the hardware TPM and the interface to manage attested transcripts. We are also able to simplify some aspects of the platform, for instance, by removing side channel protections from Tyche (e.g., cache and CPU flush). One of the implementation challenges was to adapt Tyche and the Linux driver to support large (i.e., several GBs) enclaves with our custom runtime and manage memory allocation accordingly. To enforce the integrity of the security monitor, we use the TPM and a dynamic root of trust (e.g., Intel TXT or AMD DRMT (Futral and Greene, 2013; Nie, 2007)) to implement standard measure boot (see primer in Appendix A). At the end of this process, two TPM registers (PCRs) are set with hash values that uniquely identify the security monitor binary. Later, when the security monitor needs to attest an application transcript, it will load the transcript hash in a third PCR and request a TPM Quote (a signature over select PCRs).
@@ 8.3. Custom Runtime
The execution environments provided by Argos are similar to bare-metal environments with no OS support. That means no syscalls, no program loader, no memory allocation, nor access to shared libraries. A runtime is required to support programs and modern libraries like SEAL that leverage many services from the OS. Gramine (Tsai et al., 2017) is such a runtime and is already supported by Tyche. It was conceived to securely run off-the-shelf applications on Intel SGX. As a result, it implements all Linux syscalls and also has extensive support for shared library and memory management. However, because of its versatility, it is rather large (20KLOC) and adds some non-negligible overhead (Project, 2024b). In order to reduce the TCB and improve performance, we implement our own custom runtime. It is minimal (870LOC) and consists of a simple memory allocator (Peterson and Norman, 1977) and a handful of syscall handlers required by our application. For instance, we implement our own open and read syscalls to supply randomness from RDRAND when an application attempts to read /dev/urandom. We write our runtime to interface with MUSL (musl, 2024) by interposing on system calls and providing our own handlers. That means all applications and libraries (including SEAL (Chen et al., 2017a)) need to be compiled statically and linked using our MUSL library, which requires some engineering effort to port applications but makes it possible to obtain high performance and a very light runtime.
@@ 9.1. Testbench
We prototype Argos on a local Dell Optiplex machine from 2017 equipped with an Intel Core i7-7700 processor clocked at 3.6GHz, a discrete TPM 2.0, and 8GB of RAM. We run our security monitor (adapted from Tyche (Castes et al., 2023) see Section 8.1) directly on the hardware and evaluate our benchmarks using our custom SEAL runtime or Gramine (Tsai et al., 2017). We run Ubuntu 22.04 with the Linux kernel v6.2 and SEAL 4.1 with the BFV scheme (Brakerski et al., 2014). All benchmarks are single-threaded, compiled using clang++ with the O3 optimization level.
We compare Argos with two other TEE platforms that have the following specs. SGXv2 running on an Azure server equipped with an Intel Xeon Platinum 8370C CPU with a frequency of 2.80GHz. AWS Nitro Enclaves (AWS, 2024a) running on an AWS server equipped with an Intel Xeon Platinum 8259CL CPU with a frequency of 2.50GHz. The SGXv2 setup uses Gramine as a runtime while Nitro Enclave uses a docker container environment that runs its own Linux kernel. Note that under our threat model, both SGXv2 and AWS Nitro Enclaves are insecure and are only used as comparison points for performance. See Table 7 for a more complete comparison. All numbers are obtained by averaging values over 10 runs.
@@ 9.2. TCB Evaluation
Detailed count for lines of code (LOCs) can be found in Table 3. Our TCB is small with a minimum of 40KLOC for our PIR application and a total of 60KLOC for our PSI application. Note that we did not account for SMM code, the microcode used in the DRoT, or the TPM as it is not public information, and we consider these elements as part of the hardware. This should be compared to other TEEs and system projects that try to keep a small footprint. For instance, Nitro Enclaves include an entire Linux kernel which blows up their TCB in the range of MLOC, several orders of magnitude bigger than our TCB. On the other hand, SGX would not include the SMM code or the security monitor (18KLOC), but relies on a significant amount of microcode. As a point of comparison, XEN (Barham et al., 2003) (a popular mini-hypervisor), and Coreboot (Borisov, 2009) (a minimal BIOS) each total 200KLOC. On the other hand, micro-kernels such as SeL4 (Klein et al., 2009) are in a similar size range with tens of thousands of lines of code.
@@ 9.3. Microbenchmarks
The vTPM signature requires a vmcall and a context switch to the security monitor, which incurs a small overhead (about 100s) compared to the application-level signature.
@@
However, they are only 5 faster than the discrete TPM at best.
@@
Enclave Fixed Costs. We evaluate the different fixed costs of different operations over the lifetime of an enclave. To obtain measurements, we instrument a dummy enclave for which we allocate 1GB of memory. Setup time is measured between the time a command line program requests the operating system to start an enclave and when execution reaches the main function of the enclave application. This means it includes the allocation of enclave resources, loading of the binary, page table setup, binary measurement, and runtime initialization. We can see that Nitro Enclaves have the highest setup time mostly due to the fact that they contain an entire Linux kernel that needs to be booted. SGXv2 also presents some large startup time. This is mostly due to all the special instructions required to start an enclave and the time it takes Gramine to initialize. As indicated in Gramine’s documentation, initialization can initiate more than 200 OCALLs, which all incur expensive context switches with architectural flushes (CPU caches and state) that cost 8,000 — 12,000 cycles each (Project, 2024b). This overhead can also be observed when running Argos with Gramine. In comparison, our custom runtime is initialized with minimal overhead, which might be essential for latency-sensitive applications.
@@
For Nitro Enclaves, it requires to wait for Linux to shutdown, which also adds delay.
@@
Transcript Size. Our transcript is quite small and composed of 2 PCR values used by the DRoT, the initial value of some key registers for the execution environment, the hash of our binary, the hash of messages exchanged by the application, the hash of some server input if needed. It also contains the signature and the TPM public key along with the TPM certificates endorsed by the manufacturer. Concretely, we measure a small transcript size of only 1,407 bits, negligible in front of the size of an FHE ciphertext (500KB).
@@ 9.4. FHE Evaluation
We evaluate the overhead of circuit-level vFHE in Argos compared to semi-honest FHE. Results can be seen in Table 6. We use the same benchmarks as previous work on vFHE (Viand et al., 2023) and compile and evaluate the different circuits using the SEAL library (Chen et al., 2017a). The results from Viand. et al. highlight the impractical performance overheads when using cryptographic proofs such as Bulletproofs (Bünz et al., 2018), Aurora (Ben-Sasson et al., 2019), Groth16 (Groth, 2016) or Rinocchio (Ganesh et al., 2023), to build vFHE: between 6 and 7 orders of magnitude, with the exception of Rinocchio for the Tiny benchmark. Their semi-honest baseline is also relatively slow, hinting that they might have disabled hardware acceleration. Their implementation of FHE-in-TEE also shows poor performance probably due to the limitation of SGXv1 (limited 128MB EPC memory and unreliable timers). For a fair comparison, we run the benchmarks of SGXv2 and Nitro Enclaves. We compiled the benchmarks on one machine and used the same binary across or different setups. Here, the performance looks much better with single-digit overheads and even small speedups for Nitro Enclaves. Indeed, once setup, running a program in an enclave is equivalent to running it with bare-metal performance. We tried to keep the runtime as similar as possible between the different configurations, but small differences might explain these results. For Argos, we evaluate two different configurations. One using the Gramine runtime (Argos+G) and one using our custom runtime. Our custom runtime performs better than Gramine. This might be for a couple reasons including that the MUSL library tends to be faster than Glibc, our custom runtime simply ignores interrupts sent by the OS scheduler (the enclave is not de-scheduled), and our custom syscall handlers handle less corner cases and are more efficient (for instance, to allocate memory). In conclusion, Argos shows performance comparable to commercial TEEs while offering better security. Our custom runtime makes it possible to reduce the TCB and improve performance at the cost of some engineering effort to port applications.
@@ 9.5. Authenticated Private Information Retrieval
We evaluate the overhead of authenticated PIR using Argos compared to semi-honest unauthenticated PSI. We use the SealPIR library from Angel et. al. (Angel et al., 2018). We use the BFV scheme with polynomials of degree 4096. The plaintext modulus is set to 20. We evaluate performance for a database of size where each element measures 128 bytes. We allocate 4GB of static memory to our enclave. We instrument the code to hash the database, queries, and responses and to attest the final transcript once all queries have been processed. Results can be found in Table 6. We compare two configurations for Argos. One using the Gramine runtime (Argos+G) and one using our custom runtime. For Argos+G, the server setup –which is a fixed cost– shows an overhead of mostly due to overhead setting up the enclave (+55%) and hashing the database (+16%). For query processing, we measure an overhead of for fetching one element. Here, hashing the response before sending it to the client takes a negligible amount of time. Taking into account the signing of the final transcript, the overhead goes up to for retrieving one element. Batching queries here would amortize costs down to a minimum of for numerous queries. For Argos using our custom runtime, overhead goes down to 21% for setup (mostly due to hashing of the batabase) and 1% for query processing. This is due to our more efficient handling of system calls. The only cost left is attestation that can be amortized by batching requests (see Section 6.2). Once again, our custom runtime offers better performance than Gramine. For all setups, Argos’s transcript incurs negligible (¿1%) communication overhead.
Recent work on single-server authenticated PIR using cryptographic constructions by Colombo. et al., while providing better security guarantees, show a performance overhead of 30-100 (Colombo et al., 2023). Other recent work, VeriSimplePIR (de Castro and Lee, 2024), shows more reasonable online overheads (13%-20%) but incurs significant offline communication of the same order of magnitude as the database itself (2GiB communication for a 4GiB dataset). This also makes updating the database extremely difficult, while our scheme only requires updating a Merkle tree and communicating the commitment.
@@ 9.6. Authenticated Private Set Intersection
Attested Transcript. Our remote attestation scheme makes it possible to amortize the cost of attestation over messages and queries. In the PSI scheme we use, each query requires two rounds of communication (one round for OPRF, and one for the query itself), plus one initial attestation, which would bring the overhead for one query to . Instead, we only need to pay the cost of one attestation operation per batch. The PSI scheme we use already supports batching, but only up to a certain client set size (this is a limitation of FHE-based PSI schemes). For our given server set size, we empirically measure a non-negligible failure rate for client set size above elements. A client who would want to compute the intersection of a bigger set would have to do so using several queries. That means that the maximum overhead of for a client set size of 1 can be amortized for larger client set sizes. For instance, for a client set size of , computing the complete intersection would take 10 queries and overhead would be reduced back to . This phenomenon would be heightened for more efficient PSI schemes as the cost of TPM signing is constant.